# QMUL-PH-25-01

### v1

@@ 1.2 Results
[…]
@@
where ${\cal E}_{{\rm p}}$ is a specific particular solution to (1.1) constructed explicitly in [9, 10]. Our main interest here is the second term, which is a solution to the homogeneous equation. Here $H_{\Delta}(\tau)$ denotes a particular Eichler integral (4.14) of the Hecke-normalised cusp form $\Delta\in\mathcal{S}_{2s}$ with $\mathcal{S}_{2s}$ the vector space of holomorphic cusp forms of weight $2s$. The coefficients $\lambda_{\Delta}(s;s_{1},s_{2})$ were determined in a multitude of cases in [9, 10], demonstrating that these contain the completed $L$-functions, $\Lambda(\Delta,t)$, of the corresponding holomorphic cusp form.
For the case $s_{1},s_{2}\in\mathbb{N}+1/2$ and eigenvalue $s\geq s_{1}+s_{2}+1$ of opposite parity, we use the method presented in [23] to construct a particular solution to (1.1) and show that whenever ${\rm dim}\,\mathcal{S}_{2s}\neq 0$ this does not correspond to the GES solution. In such cases, a suitable multiple of the solution of the homogeneous equation must be added to the particular solution to ensure that the complete expression is the modular invariant solution ${\cal E}(s;s_{1},s_{2};\tau)$. We show that again in this case, the GES solution to (1.1) with $s_{1},s_{2}\in\mathbb{N}+1/2$ and eigenvalue $s\geq s_{1}+s_{2}+1$ of opposite parity must take the same form (1.6). Using the results of [22] concerning certain convolutions of divisor sigma functions, we find an analytic expression for the coefficients of the solution of the homogeneous equation, which is given as
@@
Conjecture 2. The modular invariant function ${\cal E}_{i,j}^{w}(\tau)$ defined in (1.4) is a rational linear combination of generalised Eisenstein series that admits a four-dimensional lattice sum representation, which receives no contributions from $L$-values of holomorphic cusp forms.
For MGFs that are given by rational linear combinations of GESs with $s_{1},s_{2}\in\mathbb{N}$ and $s_{1},s_{2}\geq 2$ and eigenvalue $s$ with the same parity as $w=s_{1}+s_{2}$, this is indeed a conjecture since we do not prove (1.7) for the coefficient of the solution of the homogeneous equation. However, for SMFs this conjecture can be proven since these are expressed in terms of GESs with $s_{1},s_{2}\in\mathbb{N}+1/2$ and eigenvalue $s\geq s_{1}+s_{2}+1$ of opposite parity for which the proof of (1.7) follows from [22].
@@
Remarkably, precisely these particular linear combinations of GESs appear to be the only relevant ones for the study of important physical observables in superstring theory and its holographic dual CFT. This is true for MGFs for perturbative genus-one contributions to superstring amplitudes, as well as for SMFs arising in the low-energy expansion of superstring amplitudes and the large-$N$ expansion of integrated correlators in $\mathcal{N}=4$ SYM.
@@ 2.1 Theta lift: definitions and properties
In the study of superstring scattering amplitudes, being that MGFs or SMFs, we discover beautiful families of non-holomorphic modular invariant functions. The simplest examples of MGFs and SMFs are the non-holomorphic Eisenstein series, $E(s;\tau)$, with $s\in{\mathbb{N}}^{+}$ and $s\in{\mathbb{N}}+{1\over 2}$, respectively. Some properties of these modular functions, which enter into the source terms in (1.1), are reviewed in appendix A.
However, the space of modular invariant functions needed to discuss higher loop MGFs and higher order coefficients in the low-energy effective action of the scattering amplitudes in type IIB becomes significantly more complicated and goes beyond the standard world of non-holomorphic Eisenstein series or Maass functions. Importantly, both MGFs and the SMFs considered here are closely related, albeit for very different physical reasons, to a particular non-holomorphic modular function of three complex variables that can be expressed in terms of an indefinite theta-series.
@@
$[{\rm SO}(2;\mathbb{R})\times{\rm SO}(2;\mathbb{R})]\backslash{\rm SO}(2,2;\mathbb{R})/{\rm SO}(2,2;\mathbb{Z})\,,$ (2.1)
@@
[…]
@@
Before moving on to discuss the construction of the functions $A_{i,j}(\rho)$ and the properties of their theta lifts (2.10) in more detail in section 2.2 we present a different parametrisation for the indefinite theta series $\Gamma_{2,2}$, which is related to the string-theory origins of MGFs [1] and SMFs [11, 12]. One property we immediately note is that (assuming convergence) the result of performing a quadruple Poisson resummation of the integers in the theta-series (2.2) leads to $\Gamma_{2,2}(\tau;\rho,iV)=\Gamma_{2,2}(\tau;\rho,iV^{-1})$ so that by changing variables $V\to 1/V$ in (2.10) we find that ${\cal E}_{i,j}^{w}(\tau)$ satisfies the functional relation
@@ 2.2 Modular local polynomials and local Maass functions
For the case of interest the discriminant we want to consider is $D=1$. As we show in appendix B that the connected component $\mathcal{C}_{\rho^{*}}\subset\mathfrak{H}\setminus E_{1}$, which contains $\rho^{*}\coloneqq e^{\frac{\pi i}{3}}$, i.e. $\rho\in\mathcal{C}_{\rho^{*}}$, coincides with a fundamental domain of $\Gamma_{0}(2)$, i.e.
@@
The function $A_{i,j}(\rho)$ in the integrand of the theta lift (2.10) is modular invariant, so it has zero holomorphic and anti-holomorphic weight with respect to all $\gamma\in\Gamma_{0}(2)$. It can be constructed by successive applications of the Maass raising operator $D_{k}$ to a modular polynomial $\mathcal{P}(\rho)$ of even negative weight, where $D_{k}\coloneqq\partial_{\rho}+k/(\rho-\bar{\rho})$ maps a modular form of weight $(k,0)$ to a modular form of weight $(k+2,0)$. More explicitly, starting with a modular form of weight $-2n$ we can raise it to a real-analytic modular function by acting with the iterated Maass derivative operator
@@
where $(x)_{m}\coloneqq x(x+1)\dots(x+m-1)$ is the Pochhammer symbol. We note that the Maass operator $D_{k}$ satisfies the property $\Delta_{k+2}\cdot D_{k}-D_{k}\Delta_{k}=-kD_{k}$, where $\Delta_{k}\coloneqq 4D_{k-2}\,\rho_{2}^{2}\partial_{\bar{\rho}}$ is the Laplacian acting on weight $k$ modular forms. As a consequence, given that a modular local polynomial, $\mathcal{P}(\rho)$, of weight $k=-2n$ is a holomorphic function and therefore satisfies $\Delta_{-2n}\mathcal{P}(\rho)=0$. It follows that $A(\rho)\coloneqq D_{-2n}^{(n)}\mathcal{P}(\rho)$ is a modular function satisfying the Laplace eigenvalue equation
@@
where $i,j\in\mathbb{N}$ such that $k=-2n=-2(3i+j)$, and
@@
[…]
@@ 3 Inhomogeneous Laplace Equation
where ‘$\rm{{b.t.}}$’ denotes the boundary terms that arise from integrating the Laplacian $\Delta_{t}$ by parts.
@@ 3.1 Boundary terms
where ‘${\rm{{b.t.}}}$’ represents boundary terms.
@@
[…]
@@
×$\times\Big{[}E(\tfrac{w+(i+j-2m-2n-1)}{2};\tau)E(\tfrac{w-(i+j-2m-2n-1)}{2};\tau){-}E(w;\tau)\Big{]}$
@@
[…]
To obtain the complete boundary term we simply need to sum over the boundary terms in (3.28) arising from (3.20). After changing summation variables $(n,m)\to(n,r{\coloneqq}n{+}m-\tfrac{1}{2}(i+j-1))$ and inverting the Laplace equation (3.11) (with the caveat explained in footnote 6) we arrive at
@@ 3.2 Alternative approach
It is convenient to use the symmetries of $A_{i,j}(\rho)$ to map the original domain $\mathcal{F}_{\rho}$ into a three-fold covering of the fundamental domain of ${\rm SL}(2,\mathbb{Z})$ given by ^$\hat{{\cal F}}_{\rho}={\rm SL}(2,\mathbb{Z})\backslash\mathfrak{H}$. In that case the two vertical boundary lines in ${\cal F}_{\rho}$, $\rho_{1}=0$ and $\rho_{1}=1$, are respectively mapped to the lines $\rho_{1}=0-\epsilon$ and $\rho_{1}=0+\epsilon$ in ^$\hat{\cal F}_{\rho}$, with $\epsilon\to 0^{+}$. If we now define $\hat{A}_{i,j}(\rho)$ as the mapping of the local Maass function $A_{i,j}(\rho)$ to ^$\hat{{\cal F}}_{\rho}$, we see that the relation (3.35) implies the existence of a discontinuity in $\partial_{\rho_{1}}\hat{A}(\rho)$ at $\rho_{1}=0$. It therefore follows from $\partial^{2}_{\rho_{1}}|\rho_{1}|=2\delta(\rho_{1})$ that inside the domain ^$\hat{{\cal F}}_{\rho}$ we have
@@
A crucial consequence of this fact is that on the fundamental domain ^$\hat{{\cal F}}_{\rho}$ the Laplace equation (2.30) is modified to
@@
The contribution of the boundary terms to the Laplace equation may then be determined by an argument similar to (3.8). After changing the integration domain in the theta lift (2.32) from ${\cal F}_{\rho}$ to its triple cover ^$\hat{{\cal F}}_{\rho}$, acting with the $\tau$ Laplacian gives
[…]
@@
[…]
@@
$[\Delta_{\tau}-s(s-1)]\mathcal{E}^{w}_{i,j}(\tau)\sim 3\sum_{p_{1},p_{2}\in\Lambda^{\prime}}\int_{0}^{\infty}x^{\frac{w-2}{2}}y^{\frac{w-2}{2}}\,c_{i,j}(\sqrt{y/x})\,e^{-\pi x\frac{|p_{1}|^{2}}{\tau_{2}}-\pi y\frac{|p_{2}|^{2}}{\tau_{2}}}\,{\rm d}x\,{\rm d}y\,.$ (3.40)
In writing this expression we have assumed that the integration over $\rho_{2}$ can be extended from the range $[1,\infty)$ to the semi-infinite line $[0,\infty)$ by exploiting symmetry of the integrand under $\rho_{2}\to 1/\rho_{2}$. However, this ignores a subtlety that arises from the fact that the manifest $\mathfrak{S}_{3}$ symmetry of (2.33) has been lost in mapping the domain ${\cal F}_{\rho}$ to ^$\hat{\cal F}_{\rho}$ and the symmetry $\rho_{2}\to 1/\rho_{2}$ only exists if we remove the restriction $p_{1}+p_{2}\neq 0$. In other words, the expression (3.40) includes the $p_{1}+p_{2}=0$ term but the $p_{1}=0$ and $p_{2}=0$ terms must be omitted in order to avoid obvious divergences in the expression for the source term. So in order to maintain $\mathfrak{S}_{3}$ symmetry we must subtract the $p_{1}+p_{2}=0$ term from (3.40) even though its presence does not cause a divergence. From (3.30) it is easy to see that subtracting the $p_{1}+p_{2}=0$ term adds a term proportional to a single non-holomorphic Eisenstein series, while leaving the bilinear source term unaffected.

### v2

@@ 1.2 Results
where the eigenvalue $s=3i+j+1$ and the explicit expression of the general coefficients $b_{i,j}(w,r)$ is given by (3.54). All MGFs are obtained by considering ${\cal E}_{i,j}^{w}(\tau)$ when $s=3i+j+1$ and $w$ have the same parity with $1\leq s\leq w-2$ and $w\geq 3$. Conversely, SMFs are generated by ${\cal E}_{i,j}^{w}(\tau)$ when $s$ and $w$ have opposite parity with $s\geq w+1$ and $w\geq 3$. [^1] It is worth noting that many of our conclusions do apply to a wider range of weights $w$, for example SMFs with $s<w$. However, we will focus on modular functions with parameters satisfying these specific ranges since they are most relevant for superstring amplitudes and the integrated correlator.
@@
where ${\cal E}_{{\rm p}}$ is a specific particular solution to (1.1) constructed explicitly in [9, 10] and subject to the boundary condition that the coefficient ot $\tau^{2s}$ in the solution of the homogeneous equation correspondng to (1.1) is zero. Our main interest here is the second term in (1.6), which is a solution to the homogeneous equation that is exponentially suppressed as $\tau_{2}\to\infty$. Here $H_{\Delta}(\tau)$ denotes a particular Eichler integral (4.14) of the Hecke-normalised cusp form $\Delta\in\mathcal{S}_{2s}$ with $\mathcal{S}_{2s}$ the vector space of holomorphic cusp forms of weight $2s$. The coefficients $\lambda_{\Delta}(s;s_{1},s_{2})$ were determined in a multitude of cases in [9, 10], demonstrating that these contain the completed $L$-functions, $\Lambda(\Delta,t)$, of the corresponding holomorphic cusp form.
For the case $s_{1},s_{2}\in\mathbb{N}+1/2$ and eigenvalue $s\geq s_{1}+s_{2}+1$ of opposite parity to $w=s_{1}+s_{2}$, we use the method presented in [23] to construct a particular solution to (1.1) and show that whenever ${\rm dim}\,\mathcal{S}_{2s}\neq 0$ this does not correspond to the GES solution. In such cases, a suitable multiple of the solution of the homogeneous equation must be added to the particular solution to ensure that the complete expression is the modular invariant solution ${\cal E}(s;s_{1},s_{2};\tau)$. We show that again in this case, the GES solution to (1.1) with $s_{1},s_{2}\in\mathbb{N}+1/2$ and eigenvalue $s\geq s_{1}+s_{2}+1$ of opposite parity to $w=s_{1}+s_{2}$ must take the same form (1.6). Using the results of [22] concerning certain convolutions of divisor sigma functions, we find an analytic expression for the coefficients of the solution of the homogeneous equation, which is given as
@@
Importantly, for the case where the GESs have half-integer indices, $s_{1},s_{2}\in{\mathbb{N}}+1/2$, a theorem on convolution identities for divisor sums proved by Fedosova, Klinger-Logan and Radchenko [22] enables us to prove 1.2 in Section 4.2. For GESs with integer indices, $s_{1},s_{2}\in{\mathbb{N}}^{\geq 2}$ 1.2 remains unproven, albeit reproducing all particular results presented in [9, 10].
@@
Conjecture 2. The modular invariant function ${\cal E}_{i,j}^{w}(\tau)$ defined in (1.4) is a rational linear combination of generalised Eisenstein series that has the form of a four-dimensional lattice sum and for which the $L$-values of holomorphic cusp forms cancel.
Consequently, while there is no known general lattice-sum representation for a single GES, the modular invariant functions ${\cal E}_{i,j}^{w}(\tau)$ do provide a four-dimensional lattice sum representation of “special” rational linear combinations of GESs. Perhaps the key defining property of these special linear combinations of GESs is the absence of contributions proportional to $L$-values of holomorphic cusp forms, despite the presence of such terms in each individual GES that appears in the linear combination.
For MGFs that are given by rational linear combinations of GESs with $s_{1},s_{2}\in\mathbb{N}$ and $s_{1},s_{2}\geq 2$ and eigenvalue $s$ with the same parity as $w=s_{1}+s_{2}$, this is indeed a conjecture since we do not prove (1.7) for the coefficient of the solution of the homogeneous equation. However, for SMFs 1.2 can be proven since these are expressed in terms of GESs with $s_{1},s_{2}\in\mathbb{N}+1/2$ and eigenvalue $s\geq s_{1}+s_{2}+1$ of opposite parity to $w=s_{1}+s_{2}$ for which the proof of (1.7) follows from [22].
@@
Furthermore, in Section 4.3 we summarise a counting argument, first noticed for the MGF case in [9], which shows that any rational linear combination of GESs relevant for the discussion of MGFs and SMFs and for which the cusp form contribution cancels out must be given by a linear combination with rational coefficients of theta lifts ${\cal E}_{i,j}^{w}(\tau)$.
Remarkably, precisely these particular linear combinations of GESs appear to be the relevant ones for the study of important physical observables in superstring theory and its holographic dual CFT. This is true for MGFs for perturbative genus-one contributions to superstring amplitudes, as well as for SMFs arising in the low-energy expansion of superstring amplitudes and the large-$N$ expansion of integrated correlators in $\mathcal{N}=4$ SYM.
@@ 2.1 Theta lift: definitions and properties
The simplest examples of MGFs and SMFs are the non-holomorphic Eisenstein series, $E(s;\tau)$, with $s\in{\mathbb{N}}^{+}$ and $s\in{\mathbb{N}}+{1\over 2}$, respectively. Some properties of these modular functions, which enter into the source terms in (1.1), are reviewed in appendix A.
However, the space of modular invariant functions needed to discuss higher loop MGFs and higher order coefficients in the low-energy effective action of type IIB scattering amplitudes in becomes significantly more complicated and goes beyond the standard world of non-holomorphic Eisenstein series or Maass functions. Importantly, both MGFs and the SMFs considered here are closely related, albeit for very different physical reasons, to a particular non-holomorphic modular function of three complex variables that can be expressed in terms of an indefinite theta-series.
@@
${\rm SO}(2;\mathbb{R})\times{\rm SO}(2;\mathbb{R})]\backslash{\rm SO}(2,2;\mathbb{R})/{\rm SO}(2,2;\mathbb{Z})\,$, (2.1)
@@
${\cal E}_{i,j}^{w}(\tau)=\int_{\mathbb{R}^{+}\times\mathcal{F}_{\rho}}V^{{w}}A_{i,j}(\rho)\Gamma_{2,2}(\tau;\rho,iV)\,2\frac{{\rm d}V}{V^{2}}\frac{{\rm d}^{2}\rho}{\rho_{2}^{2}}\,$, (2.10)
where, as will appear natural later, the domain of integration $\mathcal{F}_{\rho}$ is the fundamental domain with respect to the congruence subgroup [^3] The congruence subgroup $\Gamma_{0}(2)$ is defined as $\Gamma_{0}(2)\coloneqq\{\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in{\rm SL}(2,\mathbb{Z})\,|\,c\equiv 0\,\rm{mod}\,2\}$. $\Gamma_{0}(2)$, i.e. $\mathcal{F}_{\rho}=\Gamma_{0}(2)\backslash\mathfrak{H}$. The parameter $w$ is a non-negative integer, $w\in\mathbb{N}$. We will shortly provide some string theory arguments that justify the importance of the integral representation (2.10) in the discussion of MGFs as well as in the context of higher-derivative corrections to the low-energy expansion of type IIB string theory.
sWe emphasise that (2.10) is quite an unusual type of theta-lift in that the integrand does not seem to be a modular function in the $\nu$ variable. It is possible to think of the $V$ integral as being originally defined over the fundamental domain $\mathcal{F}_{\nu}={\rm SL}(2,\mathbb{Z})\backslash\mathfrak{H}$ where the integrand $V^{w}$ is replaced by a modular invariant function $F(\nu)$. We choose this function to be defined by the Poincaré series $F(\nu)=\sum_{\gamma\in\Gamma_{\infty}\backslash{\rm SL}(2,\mathbb{Z})}\varphi(\gamma\cdot\nu)$ where $\Gamma_{\infty}=\{\pm\left(\begin{smallmatrix}1&n\\ 0&n\end{smallmatrix}\right)\,|\,n\in\mathbb{Z}\}$ is the Borel stabiliser of the cusp and where we consider the distribution-valued Poincaré seed $\varphi(\nu)\coloneqq V^{w}\delta(B)$. By performing the standard unfolding trick for $F(\nu)$ we change the domain of integration from $\mathcal{F}_{\nu}={\rm SL}(2,\mathbb{Z})\backslash\mathfrak{H}$ to the strip $\Gamma_{\infty}\backslash\mathfrak{H}$ while simultaneously replacing $F(\nu)$ by its Poincaré seed $\varphi(\nu)$. The integral over ${\rm Re}(\nu)=B$ can then be easily performed thanks to the delta function, thus showing that (2.10) is indeed a theta-lift of a distributional automorphic form [^4] We thank Kathrin Bringmann for related discussions.. With a slight abuse of standard terminology, we still refer to (2.10) as a theta-lift.
@@
Before moving on to discuss the construction of the functions $A_{i,j}(\rho)$ and the properties of their theta lifts (2.10) in more detail in section 2.2 we present a different parametrisation for the indefinite theta series $\Gamma_{2,2}$, which is related to the string-theory origins of MGFs [1] and SMFs [11, 12]. One property we immediately note is that (temporarily neglecting issues about regularisation) the result of performing a formal quadruple Poisson resummation of the integers in the theta-series (2.2) leads to $\Gamma_{2,2}(\tau;\rho,iV)=\Gamma_{2,2}(\tau;\rho,iV^{-1})$ so that by changing variables $V\to 1/V$ in (2.10) we find that ${\cal E}_{i,j}^{w}(\tau)$ satisfies the functional relation
@@ 2.2 Modular local polynomials and local Maass functions
For the case of interest the discriminant we want to consider is $D=1$. As we show in appendix B, the connected component $\mathcal{C}_{\rho^{*}}\subset\mathfrak{H}\setminus E_{1}$ which contains $\rho^{*}\coloneqq e^{\frac{\pi i}{3}}$ coincides with a fundamental domain of $\Gamma_{0}(2)$, i.e.
@@
The function $A_{i,j}(\rho)$ in the integrand of the theta lift (2.10) is modular invariant, so it has zero holomorphic and anti-holomorphic weight with respect to all $\gamma\in\Gamma_{0}(2)$. It can be constructed by successive applications of the Maass raising operator $D_{k}$ to a modular polynomial $\mathcal{P}(\rho)$ of even negative weight, where $D_{k}\coloneqq\partial_{\rho}+k/(\rho-\bar{\rho})$ maps a modular form of weight $(k,0)$ to a modular form of weight $(k+2,0)$. More explicitly, starting with a modular form of weight 2n-2n we can raise it to a real-analytic modular function by acting with the iterated Maass derivative operator
@@
where $(x)_{m}\coloneqq x(x+1)\dots(x+m-1)$ is the Pochhammer symbol. We note that the Maass operator $D_{k}$ satisfies the property $\Delta_{k+2}\cdot D_{k}-D_{k}\Delta_{k}=-kD_{k}$, where $\Delta_{k}\coloneqq 4D_{k-2}\,\rho_{2}^{2}\partial_{\bar{\rho}}$ is the Laplacian acting on weight $k$ modular forms. As a consequence, given that a modular local polynomial, $\mathcal{P}(\rho)$, of weight k=2nk=-2n is a holomorphic function and therefore satisfies $\Delta_{-2n}\mathcal{P}(\rho)=0$. It follows that $A(\rho)\coloneqq D_{-2n}^{(n)}\mathcal{P}(\rho)$ is a modular function satisfying the Laplace eigenvalue equation
@@
where $i,j\in\mathbb{N}$ such that k=2n=2(3i+j)k=-2n=-2(3i+j), and
@@
$\mathcal{E}^{w}_{i,j}(\tau)=\mathcal{E}^{2-w}_{i,j}(\tau)+\Big{(}{\mbox{non-holomorphic Eisenstein series}}\Big{)}\,.$ (2.39)
@@ 3 Inhomogeneous Laplace Equation
where $\rm{{b.t.}}$’ denotes the boundary terms that arise from integrating the Laplacian $\Delta_{t}$ by parts.
@@ 3.1 Boundary terms
where $\rm{{b.t.}}$}’ represents boundary terms.
@@
$c(\alpha,\beta,\gamma)=0\,,\qquad{\rm for}\qquad\alpha+\gamma\not\equiv(s-1)\,{\rm mod}\,2\,.$ (3.16)
@@
$\times\Big{[}E(\tfrac{w+(i+j-2m-2n-1)}{2};\tau)E(\tfrac{w-(i+j-2m-2n-1)}{2};\tau){-}E(w;\tau)\Big{]}$
@@
Note that we have substituted for the eigenvalue $s=3i+j+1$ in (3.28) and we can easily see that the indices of the bilinear in non-holomorphic Eisenstein series are such that $w\pm(i+j-1)\equiv w-s\,({\rm mod}\,2)$. Hence, we deduce that when $w$ and $s$ have the same parity the bilinear source terms is the product of two non-holomorphic Eisenstein series with integer indices. Similarly, when $w$ and $s$ have opposite parity the bilinear non-holomorphic Eisenstein series have half-integer indices. Finally, the third line gives terms proportional to the non-holomorphic Eisenstein series $E(w;\tau)$. Thus we have reproduce the claim (3.5).
To obtain the complete boundary term we simply need to sum over the boundary terms in (3.28) arising from (3.20). After changing summation variables $(n,m)\to(n,r{\coloneqq}n{+}m-\tfrac{1}{2}(i+j-1))$ and inverting the Laplace equation (3.11) (with the caveat explained in footnote 7) we arrive at
@@ 3.2 Alternative approach
It is convenient to use the symmetries of $A_{i,j}(\rho)$ to map the original domain $\mathcal{F}_{\rho}$ into a three-fold covering of the fundamental domain of ${\rm SL}(2,\mathbb{Z})$ given by $\hat{{\cal F}}_{\rho}={\rm SL}(2,\mathbb{Z})\backslash\mathfrak{H}$. In that case the two vertical boundary lines in ${\cal F}_{\rho}$, $\rho_{1}=0$ and $\rho_{1}=1$, are respectively mapped to the lines $\rho_{1}=0-\epsilon$ and $\rho_{1}=0+\epsilon$ in $\hat{\cal F}_{\rho}$, with $\epsilon\to 0^{+}$. If we now define $\hat{A}_{i,j}(\rho)$ as the mapping of the local Maass function $A_{i,j}(\rho)$ to $\hat{{\cal F}}_{\rho}$, we see that the relation (3.35) implies the existence of a discontinuity in $\partial_{\rho_{1}}\hat{A}(\rho)$ at $\rho_{1}=0$. It therefore follows from $\partial^{2}_{\rho_{1}}|\rho_{1}|=2\delta(\rho_{1})$ that inside the domain $\hat{{\cal F}}_{\rho}$ we have
@@
A crucial consequence of this fact is that on the fundamental domain $\hat{{\cal F}}_{\rho}$ the Laplace equation (2.30) is modified to
@@
The contribution of the boundary terms to the Laplace equation may then be determined by an argument similar to (3.8). After changing the integration domain in the theta lift (2.32) from ${\cal F}_{\rho}$ to its triple cover $\hat{{\cal F}}_{\rho}$, acting with the $\tau$ Laplacian gives
$\Delta_{\tau}\mathcal{E}^{w}_{i,j}(\tau)$ $=6\int_{\mathbb{R}^{+}\times\hat{{\cal F}}_{\rho}}V^{w}\hat{A}_{i,j}(\rho)\,\Delta_{\rho}\Gamma_{2,2}^{\prime}(\tau;\rho,iV)\,2\frac{{\rm d}V}{V^{2}}\frac{{\rm d}^{2}\rho}{\rho_{2}^{2}}\,$, (3.38)
@@
$[\Delta_{\tau}-s(s-1)]\mathcal{E}^{w}_{i,j}(\tau)$ $\sim 6\int_{\mathbb{R}^{+}\times[1,\infty)}V^{w}c_{i,j}(\rho_{2})\Gamma_{2,2}^{\prime}(\tau;i\rho_{2},iV)\,2\frac{{\rm d}V}{V^{2}}\frac{{\rm d}\rho_{2}}{\rho_{2}}\,$, (3.39)
@@
$\Delta_{\tau}-s(s-1)]\mathcal{E}^{w}_{i,j}(\tau)\sim 3\sum_{p_{1},p_{2}\in\Lambda^{\prime}}\int_{0}^{\infty}x^{\frac{w-2}{2}}y^{\frac{w-2}{2}}\,c_{i,j}(\sqrt{y/x})\,e^{-\pi x\frac{|p_{1}|^{2}}{\tau_{2}}-\pi y\frac{|p_{2}|^{2}}{\tau_{2}}}\,{\rm d}x\,{\rm d}y\,.$ (3.40)
In writing this expression we have assumed that the integration over $\rho_{2}$ can be extended from the range $1,\infty)$ to the semi-infinite line $0,\infty)$ by exploiting symmetry of the integrand under $\rho_{2}\to 1/\rho_{2}$. However, this ignores a subtlety that arises from the fact that the manifest $\mathfrak{S}_{3}$ symmetry of (2.33) has been lost in mapping the domain ${\cal F}_{\rho}$ to $\hat{\cal F}_{\rho}$ and the symmetry $\rho_{2}\to 1/\rho_{2}$ only exists if we remove the restriction $p_{1}+p_{2}\neq 0$. In other words, the expression (3.40) includes the $p_{1}+p_{2}=0$ term but the $p_{1}=0$ and $p_{2}=0$ terms must be omitted in order to avoid obvious divergences in the expression for the source term. So in order to maintain $\mathfrak{S}_{3}$ symmetry we must subtract the $p_{1}+p_{2}=0$ term from (3.40) even though its presence does not cause a divergence. From (3.30) it is easy to see that subtracting the $p_{1}+p_{2}=0$ term adds a term proportional to a single non-holomorphic Eisenstein series, while leaving the bilinear source term unaffected.